Consider the production function y = x2 for for x > 1.

(a)​Derive and graph the marginal product MP and average product AP curves.

(b)​If the price of the input is w = 1, find total cost C(y); also find marginal cost MC and average cost AC. Draw the latter two curves on the same graph.

(c)​If the input price is w = 1, can you derive the supply curve for this competitive firm? Why or why not?

With steps, please!
Answer #1

(a) The marginal product would be MP or MP or MP = 2x . The average product would be AP= or AP= or AP = x . The graph is as below.

MP,AP 00-

(b) The production function is y = x^2 , and the input demand would be as x^* = y^{0.5} . The cost of production would be C = wx or C = x , and the cost function would be hence C = x^* or C = y^{0.5} .

The marginal cost would be as MC = or come = N or MC = 0.5 y^{-0.5} or MC = \frac{1}{2y^{0.5}} . The average cost would be AC = \frac{C}{y} or AC = \frac{y^{0.5}}{y} or AC = \frac{1}{y^{0.5}} . The graph is as below.

(c) The supply function of a firm would be where MC=P, provided that MC>AVC, and since in this case TVC=TC, MC>AC. In this case, for w=1, MC<AVC for all y>0, meaning that for any P=MC, price will always be less than the cost per unit (average cost). The shutdown price is that price which can not cover the variable cost, and in this case for the given MC, the firm should shut down for any price. Hence, for w=1, there is no supply function.